2/3x - 2/3 = 0
Solve the equation

Answer by edjones(8007)

2/3x - 2/3 = 0
3*2x/3 - 3*2/3 = 3*0
2x-2=0
2x-2+2=0+2
2x=2
2x/2=2/2
x=1
Check:
2/3-2/3=0 true
Ed